OpenStudy (bohotness):

Three vertices of square ABCD are A(–1, –5), B(–3, –1), and C(1, 1). What are the coordinates of the fourth vertex?

OpenStudy (anonymous):

with 3 points you can make the equation of two lines: the line AB and the line BC

OpenStudy (anonymous):



OpenStudy (anonymous):

the line CD is parallel to the line BA and the line BC is parallel to the line AD

OpenStudy (anonymous):

the equation of a line can be done in two forms the 'slope intercept' and the 'point slope'

OpenStudy (bohotness):

@AlexandervonHumboldt2

OpenStudy (bohotness):

0,3

OpenStudy (bohotness):

9

OpenStudy (bohotness):

@AlexandervonHumboldt2

OpenStudy (bohotness):

@TheEdwardsFamily

OpenStudy (theedwardsfamily):

I do not know this! srry :(

OpenStudy (bohotness):

its okay can you help me tag ppl?

OpenStudy (theedwardsfamily):

sure

OpenStudy (theedwardsfamily):

@mathmath333 @Godlovesme @Destinymasha

OpenStudy (theedwardsfamily):

@pooja195

OpenStudy (bohotness):

@Directrix @Godlovesme

OpenStudy (godlovesme):

sorry idk :(

OpenStudy (bohotness):

:( itso kay help us tag?

OpenStudy (bohotness):

@pooja195

OpenStudy (bohotness):

@freckles @Hero @gigi2002 @grrrxxiii

OpenStudy (theedwardsfamily):

Its only 7 of my fans that r online

OpenStudy (theedwardsfamily):

now 9

OpenStudy (bohotness):

@Luigi0210 same here

OpenStudy (theedwardsfamily):

@haleyelizabeth2017 @dtan5457 @jordanloveangel

OpenStudy (haleyelizabeth2017):

I don't remember :/ sorry!

OpenStudy (bohotness):

@HaileyD help leae

OpenStudy (theedwardsfamily):

and at the same time im trying to help other peole! sheez, how many places can i be at once? :/

OpenStudy (bohotness):

HELP

OpenStudy (mathmath333):



OpenStudy (jordanloveangel):

WUT THE Questions

OpenStudy (bohotness):

Three vertices of square ABCD are A(–1, –5), B(–3, –1), and C(1, 1). What are the coordinates of the fourth vertex?

OpenStudy (bohotness):

-4 ,5?

OpenStudy (bohotness):

4<5

OpenStudy (anonymous):

srry not good at this

OpenStudy (mathmath333):

the idea here is to find the equation of line dc which will have slope same as line ab. and find the equation of line ad which will have the same slope as line bc then u have to solve the two equations for (x,y)

OpenStudy (bohotness):

its okay cake

OpenStudy (bohotness):

@ilovewolf i dont understand tho

OpenStudy (bohotness):

help PLESASE DX

OpenStudy (ilovewolf):

do you have a screenshot of the question? Maybe that could help me understand better >.>

OpenStudy (bohotness):

http://prntscr.com/66x4wp

OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{ y-1=\dfrac{(5-(-1))}{(-1-(-3))}(x-1)----------1 \hspace{.33em}\\~\\ y-5=\dfrac{(1-(-1))}{(1-(-3))}(x-(-1))----------2 \hspace{.33em}\\~\\ }\end{align}\) solve the 2 equations

OpenStudy (bohotness):

O.O CAN YOU EXPLAN THAT PLEASE

OpenStudy (bohotness):

@~Samson245~